|
Size: 18284
Comment:
|
← Revision 118 as of 2026-02-16 18:25:40 ⇥
Size: 18356
Comment:
|
| Deletions are marked like this. | Additions are marked like this. |
| Line 183: | Line 183: |
| 13:00-19:00: David. Moderate opacity (~0.4-0.5) so stayed on EMIR. Calibrated, successful pointing and focus on Mars, moved to nearby faint source SMA_1830+063 but wind gets too high for a pointing. Telescope parked at 1400. 1800 wind may be dropping enough to begin NIKA2 observations (tau < ~0.3). Begun observations with NIKA2 in polarimetry mode at 1840. Bad KID counts of 24/3/34 (Ar 1, 2, 3). Stable conditions (better than +/- 5 Jy/beam Ar2) calib1_scans on Uranus. | 13:00-19:00: David. Moderate opacity (~0.4-0.5) so stayed on EMIR. Calibrated, successful pointing and focus on Mars, moved to nearby faint source SMA_1830+063 but wind gets too high for a pointing. Telescope parked at 1400. 1800 wind may be dropping enough to begin NIKA2 observations (tau < ~0.3). Begun observations with NIKA2 in polarimetry mode at 1840. Bad KID counts of 24/3/34 (Ar 1, 2, 3). Stable conditions (better than +/- 5 Jy/beam Ar2) calib1_scans on Uranus. Project 128-25, good pointing on 0439+360 and launched 6th rep on L1482. |
Daily Reports (2026/02/10-17, NIKA2 run 79, Cryo run 91)
Observers on-site: Miren Muñoz (Tue.-Fri.), Maite Beltrán, David Taylor, Philippe André, Angel Bongiovanni (Fri. to Tue.)
AoD: Angel Bongiovanni
Operators: Gloria Liñán (night), Pedro Martínez (day)
Telescope crew and observers arrived safe. Thanks to German for his support. Lot of snow around, but melting down. The telescope looks quite clean. Raining. Ambient temp. around 0C. No wind.
Organization of shifts:
23:00-06:00 LT: Maite
06:00-13:00 LT: Philippe
13:00-20:00 LT: David
19:00-23:00 LT: Miren/Angel
Important
Focus offsets shall be measured on Ar1
Pointing offsets shall be measured on Ar2
Pool-Account usage: Accounting of observing times
Calibration observations like pointing, focus, and calib_1scan (the latter in case of NIKA2-Pol only, aimed at characterizing the instrumental polarization) which are conducted before and during project observations are part of the project time. Note that all proposals include such overheads in the time and sensitivity estimates. This holds for total power and polarimetry observations.
Other calibration observations which are important for all projects like beam maps, skydips, or calib_1scan for absolute flux calibration may be accounted under the pool account (nikaX-YY, where X = s,w and YY = year ; nikaw-25 in this case).
- When conditions are unstable and it is unclear whether project observations make sense at all, calibrations may be accounted to the pool account.
Focus and Beam maps accounting
Please fill the Table with the Ar1,2,3 focus values as follows. At best, make one table per day.
Focus sequences (NOTE: Ar2 is very noisy and this might influence the A2 focus value)
Scans nr.
Target
Ar2
Ar1
Ar3
Pol?
Notes
20260212s32-s37
1730-130
0.87
--
--
No
Starting point 1.0, only 3/5 valid scans
20260212s39-s43
1730-130
1.15
2.05
2.30
No
Starting point 1.9
20260212s47-s51
Mars
1.18
2.05
2.30
No
Starting point 2.05; best focus Ar2 from FWHM only
20260212s92-s96
Uranus
0.85
1.73
1.93
No
Starting point 2.0; best focus Ar2 from FWHM only
20260212s121-s125
Uranus
0.77
1.56
1.87
No
Starting point 1.73; best focus Ar2 from FWHM only
20260212s132-s136
Uranus
0.73
1.7
1.91
No
Starting point 1.6; best focus Ar2 & Ar3 from FWHM only
20260215s37-s42
Uranus
--
0.82
1.0
No
Starting point 1.0. Marginal fitting for array 2
20260215s53-s57
2251+158
--
0.73
0.9
No
Starting point 0.82. Marginal fitting for array 2
20260216s53-s57
SMA1653+397
0.32
1.37
1.69
No
Starting point 1.0, Ar1 & Ar3 from FWHM only
20260216s58-s62
SMA1653+397
0.55
1.68
1.69
No
Starting point 1.3, Ar1 & Ar3 from FWHM only
20260216s65-s69
1253-055
0.25
1.2
1.4
No
Starting point 1.7, Ar2 from FWHM only
20260216s71-s75
1253-055
0.31
0.97
1.12
Yes
Starting point 1.2, Ar2 from FWHM only
20260216s97-s101
K3-50A
0.5
1.43
1.67
Yes
Starting point 1.5; Very marginal fitting for Ar2
20260216s106-s110
Mars
0.32
1.31
1.48
Yes
Starting point 1.43; Ar2 from FWHM only
20260216s143-s1147
uranus
--
0.80
1.03
Yes
Starting point 1.0; Very marginal fitting for Ar2
Similarly, if any beam maps (total power or polarimetry) are taken:
Beam maps
Scan Nr
Target
Focus
TP (99 subscans) or Pol (95 subscans)
Notes
Tuesday, 10th February 2026
@9:30-15:30 LT: Maintenance time.
@15:30-20:00: All together. Heavy cloudy and raining. Mild wind. A primer about NIKA2 observations.
No observations - adverse weather. Forecast still being unfavourable.
Wednesday, 11th February 2026
00:00-06:00: Maite. Rain and wind all night.
06:00-13:00: Philippe. It's still raining mid-morning. A possible pause in the afternoon is provided, but the rain would come back at night:
Mild <wind>: ~10 m/s. Temperature outside is slightly above 0C.
13:00-19:00: David. Rain and wind.
19:00-23:00: Miren. Frozen mix, dense fog.
23:00-24:00: Maite. Fog and rain. No observations all day.
Thursday, 12th February 2026
00:00-06:00: Maite. Heavy rain, which stopped ca. 03h; then strong wind and frozen rain. Deicing ON at 06h. No observations.
06:00-13:00: Philippe. Wind speed is dropping a bit. Sky is clearing just before dawn. Deicing OFF (08:30). Pedro takes actions for checking taumeter box, telescope surface, chiller fans, etc. Auxiliary anemometer is installed.
Started moving the telescope at around 0900 after a long period of bad weather. Clear sky, but high opacity fluctuating around tau225 ~ 0.5. Started observing around 10:30 LT, after experiencing a few problems of oscillations around the commanded position. This was fixed after checking the ACU and NCS restart. The NIKA2 DAQ was also restarted with bad KID counts of 12, 4, 23 (Ar 1, 2 , 3).
Sky background is very high and atmosphere looks unstable despite the fantastic landscape we see from the control room. Difficult to see the effect of the amplifier replacing in the Nasmyth maps of array 2. Waiting for better conditions to do that.
Attempted to focus several times on various sources at various Az, El, but experienced high tracking deviation issues due to fairly strong wind. Eventually managed to get 2.5 focus sequences (s32-s37, s39-43, s47-51). First offset estimates are Ar1-Ar2= 0.9 mm and Ar3-Ar1= 0.25 mm. Then took several calib_1scans: Mars (s52), MWC349 (s55) and a skydip: s56. Last but not least, Philippe's luggage finally arrived to the telescope.
13:00-19:00: David. Taking over by doing a calib_1scan on NGC7027 (s59) and CRL2688 (s61). Opacity/instabilities are rising. Waiting for a while to see if conditions improve or not. At ~14h (LT), we confirm that opacity and its oscillations get worse, although the forecasts indicate the contrary. Switching to EMIR backup programs. One repetition for project 149-25. EMIR observations indicate tau~0.1, much lower than the taumeter reading- ice melting onto the taumeter? (see the Note about taumeter below). Switching back to NIKA2 at ~17h (LT). The NIKA2 DAQ was restarted with bad KID counts of 12, 3, 20 (Ar 1, 2 , 3). Switching to Uranus. Conditions rather stable despite tau 0.5-0.6. Peak to peak variations at 2mm of +/- 5 Jy/beam. Calib_1scan on Uranus, scan 98. Going to try and observe project 196-25 starting ~17:30. One rep. (6 scans) on Cl0024+1652 launched.
19:00-23:00: Miren. Taking over when the last scans on Cl0024+1652 were finishing. Switching to nikaw-25. Getting a pointing on Neptune (El. ~25 deg. and setting). Tried a focus on Neptune, but a couple of scans were corrupted due to high tracking deviations (wind speed rising again). Changing to Uranus. Several focus sequences launched with varied stability conditions. Valid ones are recorded in Table above. A calib_1scan on Uranus was obtained (scan 126). ALFORI was tried addressed to a flux calibration. Impossible to get a decent pointing, even with the faint option. High altitude clouds are coming. Instabilities, parametrised by the p2p variations, reached 100+ Jy/beam (2 mm). Waiting for a while. Situation does not improve and we move to EMIR backup project 147-25. Sky is unstable and opacity getting worse, we try several pointings.
23:00-24:00: Maite. Low clouds and very high opacity (~1.6). Have tried different pointing on Uranus without succeeding. Shifting to project test for a while. Impossible to point on Uranus at 30 degrees elevation. We are completely inside the cloud and the wind speed is high (we have got tracking problems). At 24:00, the telescope is parked. The forecast indicates that the clouds are going to thicken and the wind speed to increase. Another snow storm is expected for tomorrow morning.
Note about the taumeter:
- At the start of the observations, the tau_225GHZ values were about three times higher than forecasted. After initial focus corrections and flux calibrations with NIKA2, and considering the taumeter records, observations were switched to EMIR backup programs. EMIR-E1 data indicated unexpectedly low PWV values (1–1.5 mm), raising suspicions about the taumeter. Melting ice water was still dropping from the roof of the control room. Right now, access to the roof is limited due to heavy ice blocking the ladder.
- Main-beam/aperture flux fractions for Uranus obtained this morning (e.g. scan 98) were ~130/154% and 187/208% for Ar2 and 1, respectively. The calibration factors are not different from those in force in the previous run (#90). And the fractions for Uranus were ~90/100% in those arrays (see e.g. the calib_1scan 20260120s57). So, for any reason the extinction correction was being overestimated. A new inspection revealed that the remaining solid ice below the taumeter box was likely causing spillover effects, especially when taumeter samples at higher airmass. After removing it (at that time, the ice was a bit softer), tau_Z and fitting sigma values decreased significantly. A new calib_1scan on Uranus (scan 126) got the expected flux fractions (~100%). Therefore, taumeter records of 12/02/2015 from 00:00 to 16:30 UT are likely overestimated.
- For future heavy snowfalls, ensure that ice below the taumeter box (SW side) is cleared up to the level of the old TV antenna covering box, apart from the usual wiping of the box itself. Double-check opacity by using EMIR if necessary.
About focus offsets and the cold amplifier response of the Box C:
* The difference in focus between array 2 and array 1 appears to remain around 0.9-1 mm. And between array 1 and array 3, the offset is around 0.25-0.3 mm. All this despite the limited data and the fact that we did not always have reasonably stable conditions.
* Below, we have attached a screenshot of the Nasmyth coordinate map of array 2. The prevailing conditions did not allow us to obtain scans with a more decent sky background:
SB: Nice! It looks great! Thank you! Also, the difference in focus between Ar2 and Ar1 does not see to have changed since January, i.e. after the new full warm up. Also good news.
Friday, 13th February 2026
00:00-06:00: Maite. A new storm is coming. Snowfall episodes and very strong winds (estimated 30+ m/s). No observations.
06:00-13:00: Philippe. Adverse weather continues. No observations.
13:00-19:00: David. More adverse weather. No observations.
19:00-23:00: Angel. Light snowfall. Hail. Dense fog.
23:00-24:00: Maite. Frozen rain, strong winds. Deicing ON (02:00).
Saturday, 14th February 2026
00:00-06:00: Maite. Snow and freezing rain continue, along with high-speed winds.
06:00-13:00: Philippe. The adverse weather continues unchanged. Deicing is still active (high power; fortunately, the cost per kWh during nights and weekends corresponds to the off-peak tariff).
Preparing activities to clear the NIKA2 chillers and the area around the taumeter after the storm. We expect the telescope to be operational as soon as possible -safety first- afterward. Hope the wind permits operations. Current forecast:
13:00-19:00: David. Wind speed has been dropping, dense fog all the time, but atmospheric pressure is steeply rising (by the way, the famous Azores High regains its seasonal geographic position, at last). Tighter monitoring of the glycolated mix temperature of the NIKA2 He compressors by the Operators.
19:00-23:00: Angel. Foggy, high clouds. High winds now from NW direction. Deicing is still active.
23:00-24:00: Maite.
Sunday, 15th February 2026
00:00-06:00: Maite. Windy all night. Deicing continues.
06:00-13:00: Philippe. Clear and pretty skies. Too much windy for any critical operation outside/moving the telescope: ~15-20 m/s (OSN's anemometer came back to life). 10:00: wind speed decreases. Pedro begins snow and ice removal work. Taumeter readings recovered (tau_226GHz ~ 0.3, consistent with the forecast); auxiliary anemometer reinstalled. Deicing is progressing successfully, thanks to the combined action of the sun. 11:30: Moving the telescope slowly in Az. (0.1 deg/s) for inspecting surfaces, sub-reflector, etc. Looks quite clean. No need to use the platform, in principle. NIKA2 chiller fan #2 was unblocked and working OK. 12:30: deicing powered OFF. 13:00: Preparing setup for NIKA2.
13:00-23:00: David and Angel. Started at 1300, DAQ started in total power, bad kids array 1,2,3 respectively 15, 4, 29. Checking pointing behaviour with several scans. Checking that the homology corrections are correct. FWHM of point-like sources with best focus in array 1 showed a distorted beam. Checked the temperature differences of the backup structure and quadrupod are far from being stable, so the first focus sequences are not confident (only the last two in this session were recorded in Table above). Obtain flux calibrations on Uranus, Neptune, and secondary standards. In the meantime, we expect the telescope to stabilize. Launched @calib_1 scans on Uranus (20260215s35), Neptune (s44), MWC349 (s47), NGC 7027 (s48) and CRL 2688 (s49). Launched a skydip (s50). Going to try observations for project 137-25. Wind speed suddenly increases above 20 m/s. Coming back to project nikaw-25. Large tracking deviations (30.8% records lost) in a pointing scan for testing. Telescope parked at 17:30 (LT). Observations resumed at 2130. Switch to EMIR backup project 147-25. One map of 147-25 begun at 2215.
23:00-24:00: Maite. : Only 3 scans of the map of 147-25 done. Scan errors of 2.1% to 10.5%, due to tracking deviations greater than 3 arcsec (wind gusts). Telescope parked at 23:15 due to high winds.
Monday, 16th February 2026
00:00-06:00: Maite. The wind has dropped. tau_225GHz ~0.4. Starting again at 00:45 with EMIR project 102-24. Nearby calibrators are very weak. Despite using the faint calibrators script, need to do different pointings before obtaining a reasonable one. Problems with the tuning after 1.15h. Starting again from the beginning at 02:00. Tracking alarm set a couple of times despite no wind. 2 repetitions. At 3:30 the opacity is pretty low (~0.17), but the wind has increased and it is given some track errors. NIKA2 is more sensitivity to tracking errors, so we have decided to continue with EMIR and project 102-24 for another repetition and change to NIKA2 if the wind does not increase again. Still tracking errors and wind above 18 m/s for a small amount of time. Tracking errors does not seem to be totally correlated with wind strength. Decide to do another repetition of 102-25, tau_225 ~0.15. In total, 4 repetitions of 102-25. At 05:00, we shift to NIKA2 in total power. The opacity is tau_225GHz ~0.17. Bad KID counts of 15, 4, 15 (Ar 1, 2 , 3).
06:00-13:00: Philippe. Low opacity (< 0.2), stable conditions (better than +/- 5 Jy/beam Ar2) at start of shift --> Switching to polarimetry mode at 06:30 (LT) after focus sequence on 3C279=12053-055 (s65-s69) Bad KID counts of 13, 3, 27 (Ar 1, 2, 3). Repeat focus on 3C279 (s71-75) in polarized mode, then observed 1 sequence (s78-s80) on Serp-SMM4 for B-FUN=015-17 using 3c279 as IP calibrator (calib_1scans: s83, s89). Opacity and sky stability degraded (tau225 ~ 0.3 and stability between +/-5 and +/-10 Jy/beam Ar 2), but nevertheless observed another sequence (s92-s94) on SerpS-MMS18 for B-FUN. Took calib_1scans on Mars (s112) and 2251+158 (s115) for IP calibration, and a skydip (s116), all in polarimetry mode. Then had to switch to EMIR due to sky instability and tau225 > 0.4. Observed 3 maps of 10 min on M17 (1.5 loop) for EMIR project 150-25, then had to stop due to the wind getting stronger @ ~12:10 LT.
13:00-19:00: David. Moderate opacity (~0.4-0.5) so stayed on EMIR. Calibrated, successful pointing and focus on Mars, moved to nearby faint source SMA_1830+063 but wind gets too high for a pointing. Telescope parked at 1400. 1800 wind may be dropping enough to begin NIKA2 observations (tau < ~0.3). Begun observations with NIKA2 in polarimetry mode at 1840. Bad KID counts of 24/3/34 (Ar 1, 2, 3). Stable conditions (better than +/- 5 Jy/beam Ar2) calib1_scans on Uranus. Project 128-25, good pointing on 0439+360 and launched 6th rep on L1482.
19:00-22:00: Angel.
22:00-24:00: Project 175-25 (EMIR)
Tuesday, 17th February 2026
00:00-02:00: Project 175-25 (EMIR)
02:00-07:00: 087-21 (POLAMI)
07:00-10:30: Project 175-25 (EMIR)
10:30: Maintenance time.







